• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name H. Benaoum

    Log entry time 08:20:57 on July23,2005

    Entry number 149144

    keyword=Compton... problem with roc 5

    Owl shift has some problems with roc5 for compton. They tried to reboot it many 
    times but the problem persists. 
    
    I went to the hall and reset roc5. It's working fine now and we start taking 
    compton data.